mikewwe13:

Drag a statement or reason to each box to complete this proof. If −3(x+8)=−21, then x=−1. Statement Reason 1. _________ Given 2. −3(x+8)−/3=−21−3 ____________ 3. _________ Simplifying 4. x+8−8=7−8 ______________ 5. x=−1 Simplifying Subtraction Property of Equality Distributive Property of Equality Division Property of Equality x + 8 = 7 -3(x + 8) = -21

mikewwe13:

@Vocaloid

Vocaloid:

well we naturally start w/ -3(x + 8) = -21 as "given" because it's given not sure what's going on w/ step 2 can you check to see if it's written correctly

mikewwe13:

1 attachment
Vocaloid:

awesome any ideas? notice how they are ~dividing~ by -3 on each side

mikewwe13:

2nd is division

Vocaloid:

good, the Division Property of Equality now the third step is "simplifying" so which answer choice do you get if you take -3(x + 8)/(-3) and simplify? notice how the -3 cancels out

mikewwe13:

x + 8 = 7

Vocaloid:

awesome so that's blank 3 for blank 4 notice how we are ~subtracting~ 8 from each side so which property would this be?

mikewwe13:

subtraction property of equality

Vocaloid:

awesome so in order blank 1 = -3(x + 8) = -21 blank 2 = Division Property of Equality blank 3 = x + 8 = 7 blank 4 = Subtraction Property of Equality

mikewwe13:

is this correct

1 attachment
Vocaloid:

good

Vocaloid:

may I see what steps 1-3 are? they weren't included in the screengrabs you posted

mikewwe13:

hold on i'm sorry

Vocaloid:

well, statement 1 just re-states what is already given so "given" would be the appropriate choice step 2 goes from 3(x-4) to 3x - 12, which property is this?

mikewwe13:

distributive property

Vocaloid:

good step 3 adds 12 to both sides, which property is this?

mikewwe13:

addition property

Vocaloid:

awesome so step 4 is simplifying 3x - 12 + 12 = x + 4 + 12 which answer choice do you get after you take that expression + combine like terms?

mikewwe13:

3x = x + 16

Vocaloid:

good

Vocaloid:

for step 5 they subtract x from each side, which property is this?

mikewwe13:

subtraction property

Vocaloid:

good, and step 6 is simplifying so what do you get when you combine like terms for 3x - x = x + 16 - x?

mikewwe13:

2x = 16

Vocaloid:

awesome so step 7 divides both sides by 2, which property is this?

mikewwe13:

division property of equality

Vocaloid:

good last step, step 8 what is the result when you simplify 2x/2 = 16/2 ?

mikewwe13:

x = 8

Vocaloid:

awesome so to re-cap step 1: given 2: distributive 3: subtraction 4: 3x = x + 16 5: subtraction 6. 2x = 16 7; division 8: x = 8

Vocaloid:

wait

Vocaloid:

step 3 is addition sorry

mikewwe13:

yea cause i just started wondering

mikewwe13:

Which property justifies this statement? If 4x = 20, then x = 5. A. Substitution Property of Equality B. Subtraction Property of Equality C. Reflexive Property of Equality D. Division Property of Equality

Vocaloid:

any ideas? did they add, multiply, subtract, or divide?

mikewwe13:

IT'S D

mikewwe13:

Which property justifies this statement? x − 4 = x − 4 A. Substitution Property of Equality B. Reflexive Property of Equality C. Transitive Property of Equality D. Symmetric Property of Equality

mikewwe13:

I BELIEVE IS B

Vocaloid:

good

mikewwe13:

are they both correct ?

Vocaloid:

yes

mikewwe13:

i'll start a new question if that's ok

Vocaloid:

sure